OpenStudy (anonymous):

whats the square root of 16x^4y^7

OpenStudy (campbell_st):

this can be rewritten as \[\sqrt{16} \times \sqrt{x^4} \times \sqrt{y^6}\times \sqrt{y}\] can you simplify it from here

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

do you know any of the index laws...?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

makes it hard to do this question

OpenStudy (anonymous):

can you solve it and then show me

OpenStudy (anonymous):

hello

OpenStudy (campbell_st):

can you simplify \[\sqrt{16} = \sqrt{4^2} =?\]

OpenStudy (anonymous):

so is it 4x squared y to the fourth

OpenStudy (anonymous):

helloo

OpenStudy (campbell_st):

nope.... you just need to get the powers of x and y correct its \[4x^?y^? \sqrt{y}\]

OpenStudy (anonymous):

thats so confusing i still dont get it

OpenStudy (campbell_st):

use this to see if you can get the power of x \[\sqrt{y^6} = y^3\]

OpenStudy (anonymous):

x^2 and y^4

OpenStudy (campbell_st):

nope... \[\sqrt{16} timessqrt{x^4} \times \sqrt{y^3} \times \sqrt{y} = 4x^2y^3\sqrt{y}\]

OpenStudy (anonymous):

thats makes no sense

OpenStudy (campbell_st):

I think the problem is that you don't seem to have learnt about power laws or radical laws

OpenStudy (anonymous):

so if i have a problem like \[\sqrt{60a ^{6}b ^{5}c ^{3}}\]

OpenStudy (anonymous):

its 2a^3 b^4 c2 root 15bc

OpenStudy (campbell_st):

not quite a is correct.... the b needs to be split into an even and odd power \[b^5 = b^4 \times b\] same for c \[c^3 = c^2 \times c\] and split 60 into the product of a square number and another number \[60 = 4\times15.... = 2^2 \times 15\] so the problem can be split into parts \[\sqrt{2^2 a^6b^4c^2} \times \sqrt{15bc}\] the 1st square root just write the the same number and letters but just halve the powers the 2nd square root won't change

OpenStudy (anonymous):

so it would be 4a to the 3 b to the 4 and c 2 root 15bc

OpenStudy (campbell_st):

bits are correct \[\sqrt{2^2a^6b^4c^2} = 2a^3b^2c \]

OpenStudy (anonymous):

so its 2a3b2c root 15bc

OpenStudy (campbell_st):

thats right

OpenStudy (anonymous):

omg thanks i understand it now

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

its does alot

OpenStudy (anonymous):

i know you guys already got there but i did this for extra :P